For the rules for probability distributions, one can't be true:
A)
Each probability must be between 0 and 1
B) The probabilities must total 1
C) The events listed must be disjoint
D) the probability can be a negative number

Health care administrator assumes the probability of an error appearing on the patient's bill is 0.7. find the probability on 20 bills we identify

A) Four Errors b) Seven Errors   c) Eight Errors

which of the above has a higher probability to happen?


A company manufactures wheelchair batteries in batches of 6 and there is a 3% rate of defects. Find the mean number of defects per batch.
A) 5.8
B) 1.8
C) 0.2
D) 18
Answer #1

1)

D) the probability can be a negative number

2)

here this is binomial distribution with parameter n=20 and p=0.7

A) P(Four Errors)=P(X=4)=20C4(0.7)4(0.3)16 = 0.000005

B)P( seven Errors )=P(X=7)=20C7(0.7)7(0.3)13 = 0.001018

C)P( Eight Errors )=P(X=8)=20C8(0.7)8(0.3)12 = 0.03859

probability of 8 error is higher

3) mean number of defects =np=6*0.3 =0.18~ 0.2 ; option C